Sheridan, Inc., has issued a three-year bond that pays a coupon rate of 8.5 percent. Coupon payments are made semiannually. Given the market rate of interest of 4.2 percent, what is the market value of the bond? (Round answer to 2 decimal places, e.g. 15.25.)
Answer #1

Value of Bond = 1- (1+r)* Coupon * - MaturityValue (1 + r)

Where r is the discounting rate of a compounding period i.e. 4.2%/2 = 2.1% _ _ (Semi Annual )

And n is the no of Compounding periods i.e. 3years * 2 = 6 periods _ _ (Semi Annual )

= 105 (1+0.216, 0.021 1000 140.0216

= 237.269492891 + 882.76589763

= $ 1120.03

Market Value of the Bond is $ 1120.03

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.